OpenStudy (thomasfire55):

Solve this equation by completing the square. 3x^2 - 4x = -2

OpenStudy (acespeedfighter):

-2

OpenStudy (triciaal):

no direct answers. guide the asker to the correct solution.

OpenStudy (acespeedfighter):

no i was just....uh.....sorry ;-;

OpenStudy (sshayer):

for a quadratic you get two solutions.

OpenStudy (triciaal):

remember this (a-b)^2 = a^2 -2ab + b^2

OpenStudy (_citrus_):

im not really sure sorry :(

OpenStudy (welshfella):

3x^2 - 4x = -2 3[ x^2 - (4/3)x] = -2 3[ (x - 2/3 x)^2 - 4/9] = -2 (x - 2/3 x)^2 - 4/9] = -2/3 (x - 2/3 x)^2 = -2/3 + 4/9 = - 2/9

OpenStudy (welshfella):

No real roots 2 complex roots